# Modular degree and a conjecture of Watkins

### v2

@@ Abstract.
Key words and phrases:
Elliptic curves, Modular forms, Galois representations, Drinfeld modules
@@
Primary 11F30, 11L07; Secondary 11F52, 11F80
@@ 1. Introduction
Let $X_{0}(N)$ be the moduli space of the pairs $\{(E,C_{N})\mid C_{N}\subseteq E,~{}\mathrm{cyclic~{}subgroup~{}of~{}order}~{}N\}$. It is well known that $X_{0}(N)(\mathbb{C})$ can be realized as the Riemann surface obtained by the standard action of the congruence subgroup $\Gamma_{0}(N)$ on the upper half place $\mathbb{H}$. We know that $E$ is modular by the work of Wiles, Taylor-Wiles, et al. In other words, there exists a surjective morphism (defined over $\mathbb{Q}$) $\phi_{E}:X_{0}(N)\to E$, which is called the modular parametrization of $E$. Throughout the article, we shall assume that $\phi_{E}$ has a minimal degree sending the cusp $\infty$ of $X_{0}(N)$ to the identity of $E$. We denote the minimal degree as $m_{E}$, the modular degree of $E$. The main goal of this article is to study growth and some divisibility properties of $m_{E}$. To be more precise, in Section 4, we discuss the bounds of $m_{E}$ in terms of the conductor $N_{E}$ of $E$. It is conjectured that $m_{E}\ll N_{E}^{2+\varepsilon}$, and this is equivalent [29] to ABC conjecture. We discuss this briefly in Section 4. In the same section, we record the conjectural bound for a positive proportion of elliptic curves in the following form.
Theorem 1.
Let $S$ be any finite set of primes, then the proportion of elliptic curves that satisfy the conjectural degree bound is at least $1-\sum_{p\not\in S}\frac{1}{p^{2}}$
One of the other main focuses of this article is to study the following predicted by Watkins.
@@
$2^{\mathrm{rank}_{\mathbb{Z}}(E(\mathbb{Q})}\mid m_{E}.$
@@
It is not known whether the rank of elliptic curves are uniformly bounded, 111 In 2006, Noam Elkies discovered an elliptic curve with a rank of at least $28$, $y^{2}+$ $xy+y=x^{3}-x^{2}-20067762415575526585033208209338542750930230312178956502x+$ 3448161179503055646703298569039072037485594435931918036126600829629193944873224342934481611795030556467032985690390720374855944359319180361266008296291939448732243429 3448161179503055646703298569039072037485594435931918036126600829% 629193944873224342934481611795030556467032985690390720374855944359319180361266008296291939448732243429 At present, this is the highest known rank. For further information, please refer to A. Dujella. History of elliptic curves rank records. 2015. therefore the conjecture suggests that $\nu_{2}(m_{E})$ could be as large as possible. This is indeed confirmed by Dummigan-Krishnamoorthy [14], since $\omega(N_{E})-3$ is unbounded.
Dummigan [13] provided a heuristic proof of Watkins’s conjecture, using a certain Selmer group of the symmetric square of $E$ and its relation with the tangent space of a suitable deformation ring that captures relevant properties of the residual representation $\overline{\rho}:\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{GL}_{2}(E[2])$. In the case when $m_{E}$ is odd, Watkins’s conjecture predicts that $E(\mathbb{Q})$ is finite. By studying the Atkin-Lehner involution on elliptic curves of odd modular degrees, they showed that such curves have an even analytic rank, and there are at most two odd primes dividing the corresponding conductors. We shall briefly discuss this in Section 3.2. Gross [19] provided a way to relate $m_{E}$ with an element of the Picard group of $X_{0}(1)\pmod{p}$ when the elliptic curve has prime conductor $p$. Using this association, Kazalicki-Kohen [26] proved the conjecture for the odd modular degree case. They also showed that the conjecture is true for elliptic curves with the prime conductor of rank $1$, under a conjecture (see Conjecture 2) on the supersingular $j$-invariants of the elliptic curves over $\mathbb{F}_{p}$. We briefly discuss this in Section 2.2, Section 3.3, and record Theorem 5 as our take on the conjecture.
When we discuss even $m_{E}$, it is essential to note that Dummigain-Krishnamoorthy [14] established a lower bound on $\nu_{2}(m_{E})$ in terms of $\omega(N_{E})$. In fact, it turns out that $\nu_{2}(m_{E})\geq\omega(N_{E})-1$ for almost all elliptic curves over $\mathbb{Q}$. Using this and Bhargava-Shankar’s [6] result on the average rank bound, we record another heuristic argument in Theorem 4 towards Watkins’s conjecture, which says the following.
Theorem 2.
@@
When we allow $E$ to have a non-trivial 2-torsion $\mathbb{Q}$ point, one can show that $\mathrm{rank}_{\mathbb{Z}}(E(\mathbb{Q}))\leq\omega(N_{E})-1$, when $E$ is semistable. In particular, when Caro-Pasten proves $E[2]=\mathbb{Z}/2\mathbb{Z}$, Watkins’s conjecture in [10], when the number of non-split multiplicative reduction is either odd or $\omega(N_{E})$. Following the same approach, we prove an extension of this in Theorem 6 for the case $E[2]=(\mathbb{Z}/2\mathbb{Z})^{2}$.
@@
Given an elliptic curve $E/\mathbb{Q}$, we consider the family of quadratic twists $\{E^{(D)}\}$. Watkins’s conjecture has also been studied for such families. In the case when $E$ has a non-trivial 2-torsion $\mathbb{Q}$ point, Esparza-Lozano and Pasten showed in [15] that Watkins’s conjecture is true $\omega(D)$ is sufficiently large. The Crux of their argument lies in the relation of $m_{E}$ with the Peterson norm $||f_{E}||_{N}$, where $f_{E}\in S_{2}(\Gamma_{0}(N))$ be the newform associated to $E$. This association is due to Gross [19], which we recall in Section 2. When $E$ has a prime power conductor, Watkins’s conjecture for any quadratic twist $E^{(D)}$ is proved by Caro in [9]. Setzer’s classification [35] of such elliptic curves has proven useful, in addition to using the relation with $||f_{E}||_{N}$. In Section 3.5, we extend the study of twists for the cases when $E(\mathbb{Q})[2]$ is trivial. In particular, we record in Theorem 7 that Watkins’s conjecture is true for almost all the twists under the assumption of GRH and BSD conjectures.
@@
Section 2 discusses several identities about $m_{E}$. Historically, they served as tools to study several arithmetic properties of $m_{E}$. As already mentioned, the relation between $m_{E}$ and $||f_{E}||_{N}$ is recalled, and the consequence of the growth of $m_{E}$ is discussed. Many resources are available on the related subjects, and nothing original is guaranteed in this section. Instead, it serves the reader a quick introduction to the subject and provides a notational setup for the article. In Section 4, we briefly discuss the growth of $m_{E}$.
@@
In Section 3.7, we undertake the study of the function field analog of Watkins’s conjecture. Let $p>3$ be a prime and $k$ be a finite extension of $\mathbb{F}_{p}$. We define $A$ as the polynomial ring $k[T]$ and its field of fractions as $K=k(T)$. Let $E$ be an elliptic curve with conductor $\mathfrak{n}_{E}=\mathfrak{n}\infty$, where $\mathfrak{n}$ be an ideal in $A$, and $\infty$ denotes the place of $K$ associated to $T^{-1}$. We further assume that $E$ is non-isotrivial, i.e., the $j$-invariant is not in $k$ and has split multiplication reduction at $\infty$. Then we have an analog of the modular parametrization, i.e., a nonconstant morphism $\phi_{E}:X_{0}(\mathfrak{n})\to E$, where $X_{0}(\mathfrak{n})$ is the Drinfeld modular curve associated to $\mathfrak{n}$, see Section 3.7 for the definition. In this scenario, Caro [8] showed that an analog of Watkins’s conjecture holds when we view $E$ over any extension $k^{\prime}(T)$ of $K$, such that $k^{\prime}$ contains all roots of the finite part of $\mathfrak{n}$. In this section, we point out an improvement over the imposed condition on $k^{\prime}$ but at the cost of a few more additional conditions on $E$. We record this remark at Theorem 10. Furthermore, drawing the analogy with Section 3.4 and Section 3.5, we discuss Watkins’s conjecture for the families of twists $\{E^{(g)}\}$, for any polynomial $g$ having even degree. We recall Caro’s results and mention some possible improvements. Furthermore, in section 4.1, we discuss the growth properties of modular degree following [30]. It turns out that the modular degree is related to the special values of the Symmetric square $L$-function associated with $E$, whose analytic properties can be studied because Grothendieck’s theory L-function allows one to write the $L$-function as a quotient of two polynomials. We shall note that the unpleasant factor in the lower bound, which is given by the non-separable degree of the morphism $j_{E}:\mathbb{P}^{1}\to\mathbb{P}^{1}$, can be removed for a certain proportion of elliptic curves.
@@ 2.1. On the modular approach
Let $\omega_{E}$ be the invariant differential associated with $E$. Given a modular parameterization $\phi_{E}:X_{0}(N)\to E$, one can write $\phi^{*}_{E}(\omega_{E})=2\pi ic_{E}f_{E}(z)dz$, for some integer $c_{E}$. This constant (adjusting sign) $c_{E}>0$ is known to be Manin’s constant. It is expected that $c_{E}=1$ for the $\Gamma_{0}(N)$-optimal elliptic curve $E$. This is true when $E$ is semistable. In general, it is known [31, Theorem 1.3] that given any finite set of primes $S$, we have $c_{E}=O_{S}(1)$, for any elliptic curve $E/\mathbb{Q}$ that is semistable outside $S$, i.e., $p^{2}\mid N_{E}\implies p\in S$.
@@
The analytic symmetric-square $L$-function defined to be $L^{A}(Sym^{2}E,s)=\prod_{p}L_{p}^{A}(Sym^{2}E,s),$ where $L_{p}^{A}(E,s)=(1-\alpha^{2}_{p}/p^{s})^{-1}(\alpha_{p}\beta_{p}/p^{s})^{-1}(1-\beta^{2}_{p}/p^{s})^{-1}$. We have $\alpha_{p}=\overline{\beta_{p}}$ and $\alpha_{p}+\beta_{p}$ is the p th trace of Frobenius of $E$. Additionally, if $p^{2}\mid N_{E}$ then $\alpha_{p}=\beta_{p}=0$ For the description of $\alpha_{p}$ and $\beta_{p}$, please refer to [40, Section 2]. A significant connection exists between symmetric squares and modular degree, established by Shimura [37], given by:
(2) $\frac{L^{A}(Sym^{2}E,2)}{\pi i\mathrm{Vol}(E)}=\frac{m_{E}}{Nc_{E}^{2}}.$
@@
It is worthwhile to note that the local Euler factors $L_{p}^{M}(Sym^{2}E,s)=L_{p}^{A}(Sym^{2}E,s)$ whenever $p^{2}\nmid N_{E}$. If $p^{2}$ divides the conductor then $L^{A}_{p}(Sym^{2}E,s)=1$, which implies that $L^{M}_{P}(Sym^{2}E,s)=U_{p}(s)$. From equation (2), we can derive the following expression:
(3) $\frac{L^{M}(\mathrm{Sym}^{2}E,2)}{\pi i\mathrm{Vol}(E)}=\frac{m_{E}}{Nc_{E}^{2}}\prod\limits_{p^{2}\mid N}U_{p}(1).$
@@ 2.2. Gross’s association and supersingular zeroes
In this section, we assume that $N:=p$ is a prime. Denote $f_{E}$ as the associated newform of weight $2$ and level $p$. Gross in [20] associated $f_{E}$ to an element $v_{E}$ of $\mathrm{Pic}(X)$, for some suitable curve $\mathrm{X}$. The curve $\mathrm{X}$ is $X_{0}(p)\pmod{p}$, which consists of two copies of $X_{0}(1)$ that meet transversally at the points whose underlying elliptic curve is supersingular at $p$. The Picard group $\mathfrak{X}$ of $\mathrm{X}$ can be written as $\sum_{i=1}^{n}\mathbb{Z}\cdot e_{i}$, where $n=g(X_{0}(p))+1$, and each $\{e_{i}\}$ can be identified with the isomorphism class of all supersingular elliptic curves over $\overline{\mathbb{F}_{p}}$. There is $\mathbb{Z}$-linear bilinear map
@@
where the first equality is due to Mestre [28].

### v3

@@ Abstract.
Primary 11F30, 11L07
SB is supported by an institute fellowship from IISER Thiruvananthapuram and ARC grant DP230100530
SK is supported by SERB grant CRG/2023/009035.
SKP is supported by an institute fellowship from IISER Thiruvananthapuram and IMSc Chennai.
@@ 1. Introduction
Let $X_{0}(N)$ be the moduli space of the pairs $\{(E,C_{N})\mid C_{N}\subseteq E,~\mathrm{cyclic~subgroup~of~order}~N\}$. It is well known that $X_{0}(N)(\mathbb{C})$ can be realized as the Riemann surface obtained by the standard action of the congruence subgroup $\Gamma_{0}(N)$ on the upper half plane $\mathbb{H}$. We know that $E$ is modular by the work of Wiles, Taylor-Wiles, et al. In other words, there exists a surjective morphism (defined over $\mathbb{Q}$) $\phi_{E}:X_{0}(N)\to E$, which is called the modular parametrization of $E$. Throughout the article, we shall assume that $\phi_{E}$ has a minimal degree sending the cusp $\infty$ of $X_{0}(N)$ to the identity of $E$. We denote the minimal degree as $m_{E}$, the modular degree of $E$. The main goal of this article is to study growth and some divisibility properties of $m_{E}$. To be more precise, in Section 4, we discuss the bounds of $m_{E}$ in terms of the conductor $N_{E}$ of $E$. It is conjectured that $m_{E}\ll N_{E}^{2+\varepsilon}$, and this is equivalent [Murty99] to the ABC conjecture. We discuss this briefly in Section 4. In the same section, we record the conjectural bound for a positive proportion of elliptic curves. More precisely, we prove the following result.
Theorem 1.1.
Let $S$ be any finite set of primes, then the proportion of elliptic curves that satisfy the degree bound $m_{E}\ll_{S}N_{E}^{2+\varepsilon}$, is at least $1-\sum_{p\not\in S}\frac{1}{p^{2}}$.
Another main focus of this article is to examine the following prediction by Watkins.
@@
$2^{\mathrm{rank}_{\mathbb{Z}}\left(E(\mathbb{Q})\right)}\mid m_{E}.$
@@
It is not known whether the rank of elliptic curves is uniformly bounded,  [^1] In 2006, Noam Elkies discovered an elliptic curve with a rank of at least $28$, $y^{2}+$ $xy+y=x^{3}-x^{2}-20067762415575526585033208209338542750930230312178956502x+$ $34481611795030556467032985690390720374855944359319180361266008296291939448732243429$ At present, this is the highest known rank. For further information, please refer to A. Dujella. History of elliptic curves rank records. 2015. therefore the conjecture suggests that $\nu_{2}(m_{E})$ could be as large as possible. This is indeed confirmed by Dummigan-Krishnamoorthy [DK13], since $\omega(N_{E})-3$ is unbounded.
Dummigan [neil06] provided a heuristic proof of Watkins’s conjecture, using a certain Selmer group of the symmetric square of $E$ and its relation with the tangent space of a suitable deformation ring that captures relevant properties of the residual representation $\overline{\rho}:\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{GL}_{2}(E[2])$. In the case when $m_{E}$ is odd, Watkins’s conjecture predicts that $E(\mathbb{Q})$ is finite. By studying the Atkin-Lehner involution on elliptic curves of odd modular degrees, they showed that such curves have an even analytic rank, and there are at most two odd primes dividing the corresponding conductors. We shall briefly discuss this in Section 3.2. Gross [Gross85] provided a way to relate $m_{E}$ with an element of the Picard group of $X_{0}(1)\pmod{p}$ when the elliptic curve has prime conductor $p$. Using this association, Kazalicki-Kohen [KK18watkinscorrigendum] proved the conjecture for the odd modular degree case. They also showed that the conjecture is true for elliptic curves with the prime conductor of rank $1$, under a conjecture (see Conjecture 2) on the supersingular $j$-invariants of the elliptic curves over $\mathbb{F}_{p}$. We briefly discuss this in Section 2.2, Section 3.3, and record Theorem 3.4 as our take on the conjecture.
When we discuss even $m_{E}$, it is essential to note that Dummigan-Krishnamoorthy [DK13] established a lower bound on $\nu_{2}(m_{E})$ in terms of $\omega(N_{E})$. In fact, it turns out that $\nu_{2}(m_{E})\geq\omega(N_{E})-1$ for almost all elliptic curves over $\mathbb{Q}$. Using this and Bhargava-Shankar’s [BS13] result on the average rank bound, we record another heuristic argument in Theorem 3.1 towards Watkins’s conjecture, which says the following.
Theorem 1.2.
@@
When we allow $E$ to have a non-trivial 2-torsion $\mathbb{Q}$ point, one can show that $\mathrm{rank}_{\mathbb{Z}}(E(\mathbb{Q}))\leq\omega(N_{E})-1$, when $E$ is semistable. In particular, when $E(\mathbb{Q})[2]=\mathbb{Z}/2\mathbb{Z}$, Caro-Pasten proved Watkins’s conjecture in [CP22], when the number of non-split multiplicative reduction is either odd or if there is no prime of split multiplicative reduction. Following the same approach, we prove an extension of this in Theorem 3.6 for the case $E(\mathbb{Q})[2]=(\mathbb{Z}/2\mathbb{Z})^{2}$.
@@
Given an elliptic curve $E/\mathbb{Q}$, we consider the family of quadratic twists $\{E^{(D)}\}$. Watkins’s conjecture has also been studied for such families. In the case when $E$ has a non-trivial 2-torsion $\mathbb{Q}$ point, Esparza-Lozano and Pasten showed in [EP21] that Watkins’s conjecture is true when $\omega(D)$ is sufficiently large. The crux of their argument lies in the relation of $m_{E}$ with the Petersson norm $f_{E}||_{N}$, where $f_{E}\in S_{2}(\Gamma_{0}(N))$ be the newform associated to $E$. This association is due to Gross [Gross85], which we recall in Section 2. When $E$ has a prime power conductor, Watkins’s conjecture for any quadratic twist $E^{(D)}$ is proved by Caro in [caro22watkinss]. Setzer’s classification [Setzer75] of such elliptic curves has proven useful, in addition to using the relation with $f_{E}||_{N}$. In Section 3.5, we extend the study of twists for the cases when $E(\mathbb{Q})[2]$ is trivial. In particular, we record in Theorem 3.8 that Watkins’s conjecture is true for almost all the twists under the assumption of GRH and BSD conjectures.
@@
Section 2 discusses several identities about $m_{E}$. Historically, they served as tools to study several arithmetic properties of $m_{E}$. As already mentioned, the relation between $m_{E}$ and $f_{E}||_{N}$ is recalled, and the consequence of the growth of $m_{E}$ is discussed. Many resources are available on the related subjects, and nothing original is guaranteed in this section. Instead, it serves the reader a quick introduction to the subject and provides a notational setup for the article. In Section 4, we briefly discuss the growth of $m_{E}$.
@@
In Section 3.7, we undertake the study of the function field analog of Watkins’s conjecture. Let $p>3$ be a prime and $k$ be a finite extension of $\mathbb{F}_{p}$. We define $A$ as the polynomial ring $k[T]$ and its field of fractions as $K=k(T)$. Let $E$ be an elliptic curve with conductor $\mathfrak{n}_{E}=\mathfrak{n}\infty$, where $\mathfrak{n}$ be an ideal in $A$, and $\infty$ denotes the place of $K$ associated to $T^{-1}$. We further assume that $E$ is non-isotrivial, i.e., the $j$-invariant is not in $k$ and has split multiplication reduction at $\infty$. Then we have an analog of the modular parametrization, i.e., a nonconstant morphism $\phi_{E}:X_{0}(\mathfrak{n})\to E$, where $X_{0}(\mathfrak{n})$ is the Drinfeld modular curve associated to $\mathfrak{n}$, see Section 3.7 for the definition. In this scenario, Caro [caro2022watkins] showed that an analog of Watkins’s conjecture holds when we view $E$ over any extension $k^{\prime}(T)$ of $K$, such that $k^{\prime}$ contains all roots of the finite part of $\mathfrak{n}$. In this section, we point out an improvement over the imposed condition on $k^{\prime}$ but at the cost of a few more additional conditions on $E$. We record this remark at Theorem 3.12. Furthermore, drawing the analogy with Section 3.4 and Section 3.5, we discuss Watkins’s conjecture for the families of twists $\{E^{(g)}\}$, for any polynomial $g$ having even degree. We recall Caro’s results and mention some possible improvements. Furthermore, in section 4.1, we discuss the growth properties of modular degree following [PM2002]. It turns out that the modular degree is related to the special values of the Symmetric square $L$-function associated with $E$, whose analytic properties can be studied because Grothendieck’s theory L-function allows one to write the $L$-function as a quotient of two polynomials. We shall note that the unpleasant factor in the lower bound, which is given by the non-separable degree of the morphism $j_{E}:\mathbb{P}^{1}\to\mathbb{P}^{1}$, can be removed for a certain proportion of elliptic curves.
@@ 2.1. On the modular approach
Let $\omega_{E}$ be the invariant differential associated with $E$. Given a modular parameterization $\phi_{E}:X_{0}(N)\to E$, one can write $\phi^{*}_{E}(\omega_{E})=2\pi ic_{E}f_{E}(z)dz$, for some integer $c_{E}$. This constant (adjusting sign) $c_{E}>0$ is known to be Manin’s constant. It is expected that $c_{E}=1$ for the $\Gamma_{0}(N)$-optimal elliptic curve $E$. This is true when $E$ is semistable. In general, it is known [pasten2018shimura, Theorem 1.1] that given any finite set of primes $S$, we have $c_{E}=O_{S}(1)$, for any elliptic curve $E/\mathbb{Q}$ that is semistable outside $S$, i.e., $p^{2}\mid N_{E}\implies p\in S$.
Now let $f$ be any cuspform of level $N$. Then the Petersson norm $||\cdot||$ is defined by $f||_{N}=(\int_{Y_{0}(N)}|f(z)|dxdy)^{1/2}$. This can be computed explicitly by Rankin’s method, which is given by
It is then due to Deligne [Delinge85] that
@@
The analytic symmetric-square $L$-function defined to be $L^{A}(Sym^{2}E,s)=\prod_{p}L_{p}^{A}(Sym^{2}E,s),$ where $L_{p}^{A}(E,s)=(1-\alpha^{2}_{p}/p^{s})^{-1}(\alpha_{p}\beta_{p}/p^{s})^{-1}(1-\beta^{2}_{p}/p^{s})^{-1}$. We have $\alpha_{p}=\overline{\beta_{p}}$ and $\alpha_{p}+\beta_{p}$ is the p th trace of Frobenius of $E$. Additionally, if $p^{2}\mid N_{E}$ then $\alpha_{p}=\beta_{p}=0$ . For the description of $\alpha_{p}$ and $\beta_{p}$, please refer to [MW02, Section 2]. A significant connection exists between symmetric squares and modular degree, established by Shimura [shimura76], given by:
(2.2) $\frac{L^{A}(Sym^{2}E,2)}{\pi i\mathrm{Vol}(E)}=\frac{m_{E}}{Nc_{E}^{2}}.$
@@
It is worthwhile to note that the local Euler factors $L_{p}^{M}(Sym^{2}E,s)=L_{p}^{A}(Sym^{2}E,s)$ whenever $p^{2}\nmid N_{E}$. If $p^{2}$ divides the conductor then $L^{A}_{p}(Sym^{2}E,s)=1$, which implies that $L^{M}_{P}(Sym^{2}E,s)=U_{p}(s)$. From equation (2.2), we can derive the following expression:
(2.3) $\frac{L^{M}(\mathrm{Sym}^{2}E,2)}{\pi i\mathrm{Vol}(E)}=\frac{m_{E}}{Nc_{E}^{2}}\prod\limits_{p^{2}\mid N}U_{p}(1).$
@@ 2.2. Gross’s association and supersingular zeroes
In this section, we assume that $N:=p$ is a prime. Denote $f_{E}$ as the associated newform of weight $2$ and level $p$. Gross in [GK92] associated $f_{E}$ to an element $v_{E}$ of $\mathrm{Pic}(X)$, for some suitable curve $\mathrm{X}$. The curve $\mathrm{X}$ is $X_{0}(p)\pmod{p}$, which consists of two copies of $X_{0}(1)$ that meet transversally at the points whose underlying elliptic curve is supersingular at $p$. The Picard group $\mathfrak{X}$ of $\mathrm{X}$ can be written as $\sum_{i=1}^{n}\mathbb{Z}\cdot e_{i}$, where $n=g(X_{0}(p))+1$, and each $\{e_{i}\}$ can be identified with the isomorphism class of all supersingular elliptic curves over $\overline{\mathbb{F}_{p}}$. There is $\mathbb{Z}$-linear bilinear map
@@
where the first equality is due to Mestre [Mestre86].